# Auto-Adaptive PINNs with Applications to Phase Transitions

### v3

@@ 1. Introduction
The study of Physics-Informed Neural Networks (PINNs) has grown rapidly in the past several years [review, review2]. One key issue in the study of these objects is the difference between applications to stationary and time-dependent problems. Despite theoretical similarities, consistent methods for the training of PINNs on time dependent problems remain elusive. We propose that choosing efficient and intelligent sampling distributions are potentially a key to alleviating these issues.
In the simulation of statics with PINNs it is a common technique to oversample, undersample, or otherwise highlight in regions which are known to be problematic such as regions containing shocks [Shock1, Shock2], difficult boundary conditions [BoundaryPINN], or other irregularities [Javi]. This is similar in spirit to multigrid schemes used extensively in traditional finite element methods, reviewed in [multiscalerev]. In this context it is well understood that choosing a proper mesh can make the difference between a problem being completely infeasible and easily solvable. Generally in static PINNs, problem regions are identified post-hoc and then manually sampled more heavily. This allows for the gradual improvement of the static function over the network training. As the training continues it is paused periodically, at which point the sampling distributions can be edited as needed. We refer to this methodology as post-hoc sampling, as it done manually after some amount of training or detailed analysis of the particular approximation is done.
This post-hoc sampling is built on the fact that it is common in complex problems that some regions in the domain are more difficult to simulate than others. Simpler regions may have well-bounded error even without exceptionally low residual value, while other regions can explode in error with even the slightest inaccuracy. This is known as the conditioning of a problem [NLA, me]. If small residual error leads to small actual accuracy, the problem is called well-conditioned. If a small residual does not necessarily result in small accuracy, the problem is called ill or poorly conditioned. More precisely, one can often bound the actual accuracy by some power of the loss times a constant. This constant is called the condition number. If it is very large, then the problem is poorly conditioned.
@@
To alleviate these issues, several methods have been experimented with. The first of these is Extended PINNs or XPINNs, first presented in [xPINN]. This method involves the decomposition of the domain into non-overlapping subdomains. A separate PINN is then trained on each of these subdomains with an additional interior boundary condition guaranteeing that the PINNs match within the domain. Additional work has been done with this method to extend to variational problems [varxPINN] and to determine which problems the method is effective in solving [whenxPINN]. In a similar spirit [WZ] proposes the partitioning of the time variable in their time slicing II approach, which will be a focus of our study.
@@
Another method designed to address these issues is the residual adaptive (or loss adaptive) PINN. Originally introduced in [orig_resadapt], this method is explored thoroughly in [resadapt]. This involves the training of a PINN by constantly resampling according to the pointwise residual of the network. The precise nature of how this sampling is done varies, but the key idea is that the loss is reduced more effectively by highlighting regions of greater loss by sampling them more heavily. This method is extremely widely applied, but has the theoretical shortcoming that it assumes a uniform loss is optimal. For problems which are highly spatially irregular, or have regions where error grows at vastly different rates, these problems become more obvious. Of particular interest here is that the method is easily applied without prior knowledge of where the problematic regions will be, a strength that we hope to maintain in our own methodology.
@@
We propose a training method aimed at alleviating the described issues by adaptively changing the sampling distribution used in training. For our study we focus on adaptively sampling in space, though this methodology could additionally be extended to allow resampling in time. In particular we propose adaptive sampling according to arbitrary densities dependent on the network, i.e. the approximated PDE solution, and its derivatives. To do this we employ a Metropolis-Hastings sampling in parallel with the network training. This functionally allows for an ideal sampling distribution to be learned by training. Importantly this type of sampling can also be done without manual intervention during the training process, instead automatically adjusting the sampling distribution according to a pre-programmed heuristic. This allows for the sampling distribution to gradually follow problematic regions as they form, move, and change in the entire domain of the problem.
@@
Our test case for these methods is the Allen-Cahn equation, which describes the phase separation of two fluids over time [AC]. It does this through the evolution of the order parameter, which indicates the relative concentration of the two mixed fluids. We discuss these equations in more detail in section 1.2.2. In particular we experiment with an energy adaptive variant of the described auto adaptive sampling, where we sample in proportion to the pointwise energy density of the Allen-Cahn equation. As we show heuristically in 1.2.3, regions of high energy often correspond well with the problematic regions described earlier. These regions contain behavior where a small residual can invite a disproportionate amount of error. Additionally these regions are often areas more important to the continued accuracy as time is evolved.
@@
This setting highlights the issues that we hope to address with our adaptive methods. Allen-Cahn displays a separation of time and spatial scales as the interfaces quickly form and slowly move. Though these interfaces only partially describe where problematic regions occur. In particular, we contrast our results with those of [WZ], who also experiment with Allen-Cahn and the related Cahn-Hilliard model. The authors pioneer the time-slicing method by testing on the Allen-Cahn and Cahn-Hilliard systems. In their experiments, they encounter several issues that we hope to address. First, they observe consistent large error on the interfaces formed in the Allen-Cahn evolution as well as other regions where the function is away from $\pm 1$. We will show that the energy adaptive approach greatly alleviates this issue. Additionally for problems which have particularly large separation of scales, their “time-adaptive I” approach fails and they are forced to instead train several networks on the same problem in what they call the “Time-Adaptive II” approach. We find this solution undesirable, as the training of multiple networks may be expensive.
@@ 1.1. Physics Informed Neural Networks
First introduced in [PINN], a Physics-Informed Neural Network is a method of approximating a partial differential equation (PDE) using a Neural Network structure with artificial ‘data’ derived from the PDE. In this framework, we define a neural network which is itself an approximation to the solution to a PDE. We then iteratively update this approximation through gradient descent on a loss function, which measures the fidelity of the approximation to the governing equations.
@@
(1.1) $\begin{cases}{}LHS[u]=RHS(x)\text{ in }\Omega\times(0,T)\\ u(0,x)=u_{0}(x)\text{ in }\Omega\\ u(x,t)=g(x,t)\text{ on }\partial\Omega\times(0,T),\end{cases}$
@@
Here $\Omega\subseteq\mathbb{R}^{d}$ and $LHS$ represents the left hand side of the PDE, a differential operator on $u$. Meanwhile $RHS(x,t)$ represents the right hand side of the PDE, a forcing term dependent only on $x$ and $t$. The second equation in 1.1 represents the initial condition and the third equation represents the boundary condition. For simplicity we write a Dirichlet type boundary condition but Neumann, mixed, or any other boundary condition can be represented similarly.
Meanwhile in the Loss Equation 1.2, $u_{\theta}(x,t)$ is the network itself, which depends on hidden parameters $\theta$ in a composite nonlinear way and takes inputs in $\Omega\times[0,T)$. The norms represent appropriate norms on the function spaces, which can vary depending on the problem. Typically, these norms are chosen to be $L^{2}$ norms on their respective function spaces. This yields an approximate loss function which is the mean square error of the residuals on the domain. These norms are then approximated via Monte-Carlo methods during the gradient descent process, yielding an approximate loss function that is used for computation. This Monte-Carlo Sampling will become essential for our adaptive methods, as we will substitute it for other more complex sampling distributions. This amounts to switching the usual $L^{2}$ norm to a weighted $L^{2}$ norm with weights dependent on the function itself.
@@ 1.2.2. Allen-Cahn Equation
(1.6) $E[u]=\gamma_{1}\int_{\Omega}|\nabla u|^{2}dx+\gamma_{2}\int_{\Omega}\Psi(u)dx$
@@
(1.7) $\partial_{t}u=\gamma_{1}\Delta u-\gamma_{2}\Psi^{\prime}(u)$
@@
(1.8) $\Psi(s)=\frac{1}{4}(s^{2}-1)^{2}.$
This has been shown as an appropriate choice of $\Psi$, though sometimes truncations are used to avoid numerical difficulties [GinzLandApprox]. This yields the following final equation, which amount to a full description of our problem.
(1.9) $u_{t}=\gamma_{1}\Delta u-\gamma_{2}(u^{3}-u)$
@@ 1.2.3. Error Heuristic for Allen-Cahn
In the case of Allen-Cahn, we develop a heuristic to identify the high error growth regions for PINN simulation. This heuristic will be important for this implementation of Auto-Adaptive Sampling. We will study the growth of the error of a trained neural network $u_{\theta}$.
@@
(1.10) $\delta u:=\delta_{t}u_{\theta}-\gamma_{1}\Delta u_{\theta}+\gamma_{2}\Psi^{\prime}(u_{\theta})$
Then given $u$ an exact solution to (1.7) we compute the PDE governing error growth as
(1.11) $\delta_{t}\epsilon=\gamma_{1}\Delta\epsilon-(\Psi^{\prime}(u)-\Psi^{\prime}(u_{\theta}))+\delta u.$
@@
(1.12) $\delta_{t}\epsilon=\gamma_{1}\Delta\epsilon-\Psi^{\prime\prime}(u)\epsilon+\delta u.$
@@
Since we assume $\gamma_{1}<<\gamma_{2}$ and $\delta u$ is small, the $\Psi$ term (1.8) dominates the energy (1.6). A low energy region is thus one where the value stays close to $\pm 1$ while a high energy region is one with values that stay close to 0. If $u\approx\pm 1$ then $\Psi^{\prime\prime}(u)\approx-1$, and thus the corresponding term in (1.12) becomes a dampening factor. Meanwhile in the high energy regions where $u\approx 0$ we see that $\Psi^{\prime\prime}(u)\approx 2$, and the term corresponding to $\Psi^{\prime\prime}$ in (1.12) becomes an amplification factor. Thus we can see that regions of low energy naturally dampen the growth of the error while regions of high energy naturally amplify the growth of the error.
This heuristic indicates that the problematic regions in the simulation of Allen-Cahn will be areas of high energy.
@@ 2. Auto-Adaptive Sampling
In order to address the issues of moving problematic regions in time dependent PINNs, we introduce the Auto-Adaptive Sampling method. The premise of this method is to use a heuristic, possibly dependent on the network approximation itself, which indicates regions of high error growth. By sampling in proportion to this heuristic, we can directly reduce errors in the regions which are most susceptible and reach an optimal distribution of the residual loss.
In classical finite element of finite difference schemes, this can be seen as analogous to a multigrid method which refines its grid in the regions which are known to be problematic, a fairly common technique for these types of systems. However we additionally weight each point equally in the training process, leading to increased emphasis on the regions with many sample points.
This comparison is also why we decide to sample in proportion to the heuristic instead of use it as a weight. In low regularity phenomenon, the including of many sample points in the low regularity region is important to accurately capture the behavior. Using a smaller number of points with increased weight does increase the emphasis of that region in training, but may fail to accurately capture the interior behavior of the region. The use of a sampling distribution is also good for computational efficiently, which will be discussed in more detail with the Metropolis-Hastings Algorithm in section 2.1.
@@
Stated precisely, we replace the analytic formulation of the first term of 1.2 with a term of the following form
(2.1) $L_{PDE}(\theta)=\int_{0}^{T}\int_{\Omega}|LHS[u_{\theta}(x,t)]-RHS(x)|^{2}\rho(u_{\theta})dxdt.$
@@
Here $\rho(x)$ is taken proportionally to a constant $C$ plus the known heuristic function with $\int_{\Omega}\rho dx=1$ and $\rho>0$ so that $\rho$ may be interpreted as a probability density. The addition of $C$ is important, as the function must be strictly positive to be a sensible sampling distribution (there is no region which can be neglected entirely in training).
2.1. Metropolis-Hastings Algorithm
In order to efficiently sample from a probability distribution dependent on the network and its derivatives, we use the Metropolis-Hastings Algorithm [Metropolis, Hastings]. This is a sampling method which iteratively improves a collection of randomly chosen points to match the distribution given by a known probability density function. This method avoids the need to invert a distribution, and is thus ideal for settings where direct sampling is difficult.
(2.2) $\alpha=\min\left(1,\frac{f(x^{\prime})g(x^{\prime}|x)}{f(x)g(x|x^{\prime})}\right)$
It is worth noting that there is a brief history of the use of Metropolis-Hastings and the broader class of Markov-Chain-Monte-Carlo (MCMC) methods in PINNs. In particular, the results of [resadapt] suggest that this is the most effective method for the implementation of residual adaptive sampling. Though dropout methods remain more popular in residual adaptive methods, such as was employed in [WZ], because of their ease of implementation. Additionally, the Bayesian-PINN (B-PINN) [BPINN] uses similar MCMC-based inference to capture uncertainty in the network parameters, addressing noise through probabilistic modeling rather than deterministic regularization.
The Metropolis-Hastings is additionally very well-studied and lends itself very naturally to the context of Physics-Informed Neural Networks. As an iteration based method, it can be run easily in conjunction with training. Additional improvements such as multiple proposal [MultiMCMC] can nicely address common issues in PDEs and in shocks in particular. Additionally, the method is nicely parrallelizable and can be run very efficiently on GPUs [ParMCMC]. These benefits make it an exceptionally good fit for sampling in the context of Neural Networks.
As discussed below equation (2.1), it is additionally necessary to sample from a uniform distribution (as is standard for PINNs), since regions of zero energy are still important to the accuracy of our simulations. We denote by $\lambda$ the ratio of points sampled adaptively:
@@ 2.1. Metropolis-Hastings Algorithm
(2.4) $Loss_{PDE}=\lambda L_{adaptive}+(1-\lambda)L_{uniform}.$

### v4

@@ Abstract.
Key words and phrases:
Physics-Informed Neural Networks, Adaptive Sampling, Phase Transition, Allen-Cahn, Multiscale
@@ 1. Introduction
The study of Physics-Informed Neural Networks (PINNs) has grown rapidly in the past several years [1, 2]. One key issue in the study of these objects is the difference between applications to stationary and time-dependent problems. Despite theoretical similarities, consistent methods for the training of PINNs on time dependent problems remain elusive. We propose that choosing efficient and intelligent sampling distributions are potentially a key to alleviating these issues.
In the simulation of statics with PINNs it is a common technique to oversample, undersample, or otherwise highlight in regions which are known to be problematic such as regions containing shocks [3, 4], difficult boundary conditions [5], or other irregularities [6]. This is similar in spirit to adaptive mesh refinement schemes used extensively in traditional finite element methods, reviewed in [7, 8]. In this context it is well understood that choosing a proper mesh can make the difference between a problem being completely infeasible and easily solvable. Generally in static PINNs, problem regions are identified post-hoc and then manually sampled more heavily. This allows for the gradual improvement of the static function over the network training. As the training continues it is paused periodically, at which point the sampling distributions can be edited as needed. We refer to this methodology as post-hoc sampling, as it is done manually after some amount of training or detailed analysis of the particular approximation is done.
This post-hoc sampling is built on the fact that it is common in complex problems that some regions in the domain are more difficult to simulate than others. Simpler regions may have well-bounded error even without exceptionally low residual value, while other regions can explode in error with even the slightest inaccuracy. This is known as the conditioning of a problem [11, 10]. If small residual error leads to small actual accuracy, the problem is called well-conditioned. If a small residual does not necessarily result in small accuracy, the problem is called ill or poorly conditioned. More precisely, one can often bound the actual accuracy by some power of the loss times a constant. This constant is called the condition number. If it is very large, then the problem is poorly conditioned.
@@
To alleviate these issues, several methods have been experimented with. The first of these is Extended PINNs or XPINNs, first presented in [12]. This method involves the decomposition of the domain into non-overlapping subdomains. A separate PINN is then trained on each of these subdomains with an additional interior boundary condition guaranteeing that the PINNs match within the domain. Additional work has been done with this method to extend to variational problems [13] and to determine which problems the method is effective in solving [14]. In a similar spirit [15] proposes the partitioning of the time variable in their time slicing II approach, which will be a focus of our study.
@@
These ideas have motivated the broad studying of adaptive sampling methods, where training is enhanced through carefully examining the choice of sample points in network training. The first of these methods is known as residual adaptive (or loss adaptive) sampling, first introduced in [16], this method is explored thoroughly in [17]. This involves the training of a PINN by constantly resampling according to the pointwise residual of the network. The precise nature of how this sampling is done varies, but the key idea is that the loss is reduced more effectively by highlighting regions of greater loss by sampling them more heavily. This method is extremely widely applied, but has the theoretical shortcoming that it assumes a uniform loss is optimal. For problems which are highly spatially irregular, or have regions where error grows at vastly different rates, these problems become more obvious. Of particular interest here is that the method is easily applied without prior knowledge of where the problematic regions will be, a strength that we hope to maintain in our own methodology.
Additional strategies in adaptive sampling are discussed in [18, 19, 20, 21]. In particular, failure-informed PINNs (FI-PINNs) [22, 23] present a flexible adaptive sampling framework which could in principle be combined with a methodology similar to our Energy-Adaptive PINNs. However, sampling strategies based on distributions which are not residual-based are in general underexplored in the literature.
@@
We propose a training method aimed at alleviating the described issues by adaptively changing the sampling distribution used in training based not on the residual but on other network-dependent heuristics which indicate regions of high error and high error growth. For our study we focus on adaptively sampling in space, though this methodology could additionally be extended to allow resampling in time. In particular we propose adaptive sampling according to arbitrary densities dependent on the network, i.e. the approximated PDE solution, and its derivatives. To do this we employ a Metropolis-Hastings sampling in parallel with the network training. This functionally allows for an ideal sampling distribution to be learned by training. Importantly this type of sampling can also be done without manual intervention during the training process, instead automatically adjusting the sampling distribution according to the pre-programmed heuristic. This allows for the sampling distribution to gradually follow problematic regions as they form, move, and change in the entire domain of the problem. We call this method Auto-Adaptive both because of its self-referential nature and self-sufficiency once training begins. We discuss the particulars of this method in Section 2.
@@
Our test case for these methods is the Allen-Cahn equation, which describes the phase separation of two fluids over time [24]. It does this through the evolution of the order parameter, which indicates the relative concentration of the two mixed fluids. We discuss these equations in more detail in section 1.2.2. In particular we experiment with an energy adaptive variant of the described auto adaptive sampling, where we sample in proportion to the pointwise energy density of the Allen-Cahn equation. As we show heuristically in 1.2.3, regions of high energy often correspond well with the problematic regions described earlier. These regions contain behavior where a small residual can invite a disproportionate amount of error. Additionally these regions are often areas more important to the continued accuracy as time is evolved.
@@
This setting highlights the issues that we hope to address with our adaptive methods. Allen-Cahn displays a separation of time and spatial scales as the interfaces quickly form and slowly move. However, these interfaces only partially describe where problematic regions occur. In particular, we contrast our results with those of [15], who also experiment with Allen-Cahn and the related Cahn-Hilliard model. The authors pioneer the time-slicing method by testing on the Allen-Cahn and Cahn-Hilliard systems. In their experiments, they encounter several issues that we hope to address. First, they observe consistent large error on the interfaces formed in the Allen-Cahn evolution as well as other regions where the function is away from $\pm 1$. We will show that the energy adaptive approach greatly alleviates this issue. Additionally for problems which have particularly large separation of scales, their “time-adaptive I” approach fails and they are forced to instead train several networks on the same problem in what they call the “Time-Adaptive II” approach. We find this solution undesirable, as the training of multiple networks may be expensive.
@@ 1.1. Physics Informed Neural Networks
First introduced in [25], a Physics-Informed Neural Network is a method of approximating a partial differential equation (PDE) using a Neural Network structure with artificial ‘data’ derived from the PDE. In this framework, we define a neural network which is itself an approximation to the solution to a PDE. We then iteratively update this approximation through gradient descent on a loss function, which measures the fidelity of the approximation to the governing equations.
@@
(1) $\begin{cases}{}LHS[u]=RHS(x)\text{ in }\Omega\times(0,T)\\ u(0,x)=u_{0}(x)\text{ in }\Omega\\ u(x,t)=g(x,t)\text{ on }\partial\Omega\times(0,T),\end{cases}$
@@
Here $\Omega\subseteq\mathbb{R}^{d}$ and $LHS$ represents the left hand side of the PDE, a differential operator on $u$. Meanwhile $RHS(x,t)$ represents the right hand side of the PDE, a forcing term dependent only on $x$ and $t$. The second equation in 1 represents the initial condition and the third equation represents the boundary condition. For simplicity we write a Dirichlet type boundary condition but Neumann, mixed, or any other boundary condition can be represented similarly.
Meanwhile in the Loss Equation (2), $u_{\theta}(x,t)$ is the network itself, which depends on hidden parameters $\theta$ in a composite nonlinear way and takes inputs in $\Omega\times[0,T)$. The norms represent appropriate norms on the function spaces, which can vary depending on the problem. Typically, these norms are chosen to be $L^{2}$ norms on their respective function spaces. This yields an approximate loss function which is the mean square error of the residuals on the domain. These norms are then approximated via Monte-Carlo methods during the gradient descent process, yielding an approximate (or empirical) loss function that is used for computation. This Monte-Carlo Sampling will become essential for our adaptive methods, as we will substitute it for other more complex sampling distributions. This will amount to switching the usual $L^{2}$ norm to a weighted $L^{2}$ norm with weights dependent on the function itself, which will be discussed in detail in section 2.
@@ 1.2.2. Allen-Cahn Equation
(6) $E[u]=\gamma_{1}\int_{\Omega}|\nabla u|^{2}dx+\gamma_{2}\int_{\Omega}\Psi(u)dx$
@@
(7) $\partial_{t}u=\gamma_{1}\Delta u-\gamma_{2}\Psi^{\prime}(u)$
@@
(8) $\Psi(s)=\frac{1}{4}(s^{2}-1)^{2}.$
This has been shown as an appropriate choice of $\Psi$, though sometimes truncations are used to avoid numerical difficulties [26]. This yields the following final equation, which amount to a full description of our problem.
(9) $u_{t}=\gamma_{1}\Delta u-\gamma_{2}(u^{3}-u)$
@@ 1.2.3. Error Heuristic for Allen-Cahn
In the case of the Allen–Cahn equation, we develop a heuristic for identifying regions of high error growth in PINN simulations, derived from well-established finite-difference and finite-element error estimation principles. This heuristic will be important for this implementation of Auto-Adaptive Sampling. We will study the growth of the error of a trained neural network $u_{\theta}$.
@@
(10) $\delta u:=\delta_{t}u_{\theta}-\gamma_{1}\Delta u_{\theta}+\gamma_{2}\Psi^{\prime}(u_{\theta})$
Then given $u$ an exact solution to (7) we compute the PDE governing error growth as
(11) $\delta_{t}\epsilon=\gamma_{1}\Delta\epsilon-\gamma_{2}(\Psi^{\prime}(u)-\Psi^{\prime}(u_{\theta}))+\delta u.$
@@
(12) $\delta_{t}\epsilon=\gamma_{1}\Delta\epsilon-\gamma_{2}\Psi^{\prime\prime}(u)\epsilon+\delta u.$
@@
Since we assume $\gamma_{1}<<\gamma_{2}$ and $\delta u$ is small, the $\Psi$ term (8) dominates the error grown equation (12). A low energy region is thus one where the value stays close to $\pm 1$ while a high energy region is one with values that stay close to 0. If $u\approx\pm 1$ then $\Psi^{\prime\prime}(u)\approx 2$, and thus the corresponding term in (12) becomes a dampening factor. Meanwhile in the high energy regions where $u\approx 0$ we see that $\Psi^{\prime\prime}(u)\approx-1$, and the term corresponding to $\Psi^{\prime\prime}$ in (12) becomes an amplification factor.
This heuristic indicates that the problematic regions in the simulation of Allen-Cahn will be areas of high energy. Thus we can see that regions of low energy naturally dampen the growth of the error while regions of high energy naturally amplify the growth of the error when the residual is reasonably small relative to the true solution.
@@ 2. Auto-Adaptive Sampling
In order to address the issues of moving problematic regions in time dependent PINNs, we introduce the Auto-Adaptive Sampling method. The premise of this method is to use a heuristic dependent on the network approximation itself, which indicates regions of high error growth, as a sampling distribution for the network training process. By sampling in proportion to this heuristic, we can directly reduce errors in the regions which are most susceptible and reach an optimal distribution of the residual loss.
@@
Stated precisely, we replace the analytic formulation of the first term of (2) with a term of the following form
(13) $L_{PDE}(\theta)=\int_{0}^{T}\int_{\Omega}|LHS[u_{\theta}(x,t)]-RHS(x)|^{2}\rho(u_{\theta})dxdt.$
@@
Here $\rho(x)$ is taken proportionally to a constant $C$ plus a known heuristic function with $\int_{\Omega}\rho dx=1$ and $\rho>0$ so that $\rho$ may be interpreted as a probability density (and any scaling can be seen as a reweighting of the PDE loss term). The addition of $C$ is important, as the function must be strictly positive to be a sensible sampling weight (there is no region which can be neglected entirely in training).
Rather than computing this weight directly, we will use it as a sampling distribution for the training process. By sampling in this manner and weighting all points equally in the loss computation, we implicitly recover the weight without explicitly computing its value. In addition, this naturally refines our distribution to be finer in difficult regions. In classical finite element schemes, this can be seen as analogous to a adaptive mesh refinement methods, which refine their mesh in the regions which are measured to be problematic.
This comparison is also why we decide to sample in proportion to the heuristic instead of use it as a directly computed weight on sample points. In low regularity phenomenon, the including of many sample points in the low regularity region is important to accurately capture the behavior. Using a smaller number of points with a directly computed weight does increase the emphasis of that region in training, but may fail to accurately capture the interior behavior of the region. The use of a sampling distribution is also beneficial for computational efficiency, which will be discussed in more detail with the Metropolis-Hastings Algorithm in section 2.2.
With this framing, the constant $C$ can be reinterpreted as denoting the proportion of uniformly sampled points as opposed to points sampled in proportion to the heuristic. Thus we denote by $\lambda$ the ratio of points sampled adaptively:
@@
(15) $Loss_{PDE}=\lambda L_{adaptive}+(1-\lambda)L_{uniform}.$